OpenStudy (anonymous):

In simplified exponential notation, the expression a²·a -3·a=

OpenStudy (anonymous):

-3 is the exponent

OpenStudy (phi):

Did you learn that a*a is written in "short-hand" as \(a^2\) and a*a*a is \( a^3\) and a*a*a*a is \(a^4\) ? in your question \[ a^2·a^{ -3}·a= ?\] you can change the order of multiplying, like this \[ a^2 \cdot a \cdot a^{-3} \] now what is a^2 ? it is the same as a*a \[ a·a·a·a^{-3} \]

OpenStudy (phi):

you should also know that \[ a^{-3} \text{ is the same as } \frac{1}{a·a·a} \] your problem is \[ \frac{a·a·a}{a·a·a} \] any idea what that is ?